# GlucOS: Security, correctness, and simplicity for automated insulin delivery

### v3

@@ Abstract.
GlucOS combines algorithmic security, driver security, and end-to-end verification to protect against malicious ML models, vulnerable pump drivers, and drastic changes in human physiology. We use formal methods to prove correctness of critical components and incorporate humans as part of our defensive strategy. Our evaluation includes both a real-world deployment with seven individuals and results from simulation to show that our techniques generalize. Our results show that GlucOS maintains safety and improves glucose control even under attack conditions. This work demonstrates the potential for secure, personalized, automated healthcare systems. Our source code is open source.
@@ 1. Introduction
Type 1 Diabetes (T1D) is a metabolic disorder where an individual’s pancreas stops producing insulin. To compensate, they inject synthetic insulin. Computer systems, called automated insulin delivery systems, exist that use subcutaneous sensors to continuously monitor glucose concentrations (glucose is the body’s primary energy source) and inject insulin automatically using an insulin pump device (Figure 1). However, insulin is a dangerous hormone as too much insulin can kill people in a matter of hours (cryer2012severe, ) and too little insulin can kill people in a matter of days (benoit2018trends, ). For automated insulin delivery systems, the key challenge is to maintain a balance: provide enough insulin to prevent dangerously high glucose levels while avoiding excessive insulin that could lead to life-threateningly low glucose levels. In this context, security means ensuring integrity for insulin dosing against malicious attacks or inadvertant errors. Several commercial (beta_bionics, ; tandem_iq, ; omnipod, ) and open-source (loop, ; openaps, ) automated insulin delivery systems exist today, but none of them consider security as a primary design constraint.
@@
Current open-source automated insulin delivery systems are complex. For example, the OpenAPS (openaps, ) core function for calculating insulin doses consists of 1192 lines of Javascript code, 63 input and configuration parameters, and 90 branch statements. All of this logic influences the ultimate dosing decision, making reasoning about its correctness, both manually and automatically, challenging. Automated insulin delivery systems that use the OpenAPS algorithm embed it into native apps using a WebView for Javascript interpretation, adding to the overall complexity and increasing the attack surface. In contrast, our implementation of this same function is 25 lines of native and formally verified Swift code. None of the open-source automated insulin delivery systems uses formal methods, can withstand attacks on components, or has mechanisms or policies for security.
In this paper, we take on the challenge of building the first trustworthy automated insulin delivery system, called GlucOS. Our design includes: (1) an architecture where we apply separation principles for isolated and simple components, (2) a novel security mechanism and policy to enable secure ML-driven insulin delivery, (3) the ability to withstand attacks on insulin pump driver software, and (4) the application of formal methods to prove correct our insulin delivery path, the most sensitive part of the system.
Despite these efforts, we identify scenarios in automated insulin delivery that require users to take action for security. Thus, we incorporate humans as part of our defensive strategy, providing them with agency and trust to take the required, corrective actions.
Our approach is distinguished by its holistic consideration of the entire problem of automated insulin delivery. We provide clear security boundaries, mechanisms and policies to mitigate attacks on vulnerable components, and formal methods to provide assurances that our implementation is correct.
Automated insulin delivery systems are different from traditional applications that the systems and security communities focus on, but they are important. Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating the impact of some or all of these vulnerabilities by virtue of their design. Automated insulin delivery systems do not have a long list of CVEs that we can reference, although there are a few (cve1, ; cve2, ; cve3, ; cve4, ). However, our position is that we need to make sure that these systems never get CVEs, because if they do, security vulnerabilities can kill people. Thus, we focus on building a new system, using security first principles and formal methods, to avoid security vulnerabilities and mitigate their damage from the start.
@@
  • Security mechanisms and policies that protect individuals from malicious ML, vulnerable pump drivers, and drastic changes in the human’s glycemic response.
@@ 2.1. Locking down the insulin delivery path
We focus our system architecture on providing security mechanisms and policies on the insulin delivery path. The insulin delivery path includes models for calculating the amount of insulin still active in the individual (insulin can take up to six hours to fully absorb), closed-loop algorithms for calculating insulin dosing, deep neural network ML to predict future metabolic states, and drivers for programming the insulin pump. Together, these components are both the most complex, and the most sensitive from a security perspective because they are responsible for insulin delivery.
@@ 2.2. Architecture
The BioKernel runs the closed-loop algorithm every five minutes as the new CGM readings come in. Once the BioKernel receives new CGM readings, it collects the current metabolic state of the individual and sends it to a predictive ML model. The predictive ML model calculates the amount of insulin to inject based on its assessment of the individual’s current metabolic state and predictions of future metabolic states. The BioKernel then issues commands to the pump to deliver the calculated amount of insulin. A pump driver converts these commands into low-level I/O to program the pump and adjust insulin delivery.
To withstand attacks, the BioKernel serves as a reference monitor (anderson1972computer, ) to enforce security policies on the predictive ML model’s pump commands and the pump driver’s low-level I/O. This reference monitor architecture ensures that we can withstand attacks from malicious predictive ML as well as malicious pump drivers without having to know the internals of how they operate.
For predictive ML, our reference monitor enforces algorithmic security, where we bound the amount of insulin calculated by ML to stay within dynamic safety bounds computed by a simple reactive safe model. For the pump driver, we interpose on the I/O path to ensure that only the pump commands issued by the BioKernel are programmed into the pump. These checks constitute our pump invariants, which, when violated switch the pump to operate in a safe mode.
Our use of formal methods focuses on the software related to delivering insulin, which is the most risky operation in any automated insulin delivery system. We use Hoare logic to prove key functions correct, define system states and transitions to handle runtime verification check failures safely, and we introduce the notion of a novel biological invariant. Our biological invariant is an end-to-end check of our theoretical calculations of how the individual’s metabolism should behave compared to what we observe in practice, where violations of this invariant suggest a potentially catastrophic failure of the system.
@@ 2.3. Threat model
The most serious health risk for people who use synthetic insulin is low glucose, or hypoglycemia. Hypoglycemia results from an overdose of insulin. Once an automated insulin delivery system has delivered insulin, there is no way to remove it from the body. If untreated, severe hypoglycemia can lead to loss of consciousness, seizure, or death in a matter of hours. Thus, our primary security objective is to reduce or eliminate hypoglycemia that results from an attack on the system. Quantitatively, our goal is to maintain the American Diabetes Association’s recommended time in hypoglycemia of less than 4% (AmericanDiabetesAssociation2021, ).
@@
As our design philosophy centers around humans, providing flexibility for people to use whatever ML they want by loading personalized models or using any pump hardware they want is important. This flexibility helps guide our focus on providing novel security mechanisms and policies for these components. Our goal is to make sure that they stay safe even if the model they pick is malicious or the pump driver they choose has security vulnerabilities. Attackers may target automated insulin delivery systems because they can directly harm the human using the system.
We assume that the Swift type system is correct and that the underlying operating system, the CGM driver and hardware, and the pump hardware we use are free of attacks and their device communications maintain integrity. Thus, we do not consider attacks from tampered CGM sensors or pump hardware in our threat model. These are important but complementary problems to our work, which focuses on the system architecture and algorithms for automated insulin delivery.
We consider attacks from malicious predictive ML algorithms. These attacks can come from trojan attack logic embedded directly in a model, or from an adversarial attack on predictive ML. For adversarial attacks, although we assume that the CGM and insulin pump remain attack free, an adversary can manipulate these entities by serving an individual who is using our system with calculated amounts of carbohydrates in their food (e.g., at a restaurant), thus influencing the resulting CGM readings and corresponding insulin commands. In our threat model, malicious predictive ML can issue arbitrary insulin doses.
We also consider malicious pump drivers. In our threat model, malicious pump drivers can issue arbitrary pump I/O to add, drop, or modify commands sent to the pump.
Successful attacks on predictive ML or pump drivers, if unchecked, can be devastating because they directly affect insulin delivery and can be used by an attacker to induce potentially life-threatening hypoglycemia.
3. Algorithmic security: Withstanding malicious ML
@@
Currently, no automated insulin delivery systems use deep neural network ML. Some current systems use statistical methods, like linear regression or model predictive control (physio2011, ; physio2019, ; lunze2013control, ; control2019, ; openaps, ). The reason for using this more traditional form of ML is sound. It provides consistent, explainable results and has a physiological basis that people can use to vet its decisions. However, all of the recent advances in ML, like image classifiers and large language models, have come from deep neural networks, which researchers have shown also work well for predicting metabolic states for use in automated insulin delivery systems (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ). Despite this promise, none of the current automated insulin delivery systems incorporate more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” (chatgpt_hallucination, ) could have lethal consequences if it delivers an inappropriate insulin dose (cryer2012severe, ; benoit2018trends, ).
In this section, we discuss the design and implementation of GlucOS’s algorithmic security mechanisms to enable predictive models, including deep neural network ML, to run safely in automated insulin delivery systems.
@@ 3.1. Mechanism fundamentals
Our key novel insight underpinning our security mechanism design is that all correct insulin delivery algorithms will deliver the same amount of insulin over a sufficiently long period. People inject insulin to facilitate glucose absorption after eating. Digestion converts food to glucose, and insulin facilitates the absorption of glucose to tissue (e.g., the brain) where the body can use it as an energy source. Since insulin dosing is driven by food consumption, all correct algorithms should dose the same amount of insulin to cover glucose from digestion.
However, the timing of this insulin delivery matters for maintaining balanced glucose levels (Section 9). Insulin activation is a slow process, with injected insulin taking up to six hours to fully activate. To meet this challenge, our algorithmic security mechanism pairs two models. Our first model is a predictive model that anticipates future metabolic states and doses insulin proactively. Our second model is a reactive safe model that measures the current metabolic state and doses insulin based only on what it can measure currently. This pairing allows us to balance proactive insulin dosing for faster responses with a more conservative approach for safety.
Architecturally, in a novel approach for automated insulin delivery systems, the predictive model runs outside of our trusted computing base while the reactive safe model is part of the trusted computing base. This separation enables individuals to use whatever predictive models they want, while maintaining security. It also defends against malicious predictive models, as all insulin delivery decisions are ultimately vetted by the reactive safe model. This design also enables people to update their predictive models without violating core safety guarentees.
Conceptually, the predictive model is the primary model in our system and should control the pump most of the time. We use the reactive safe model for accounting so that we can track how far the predictive model deviates from a known-to-be-safe baseline. Then, we bound the size of this deviation. By combining these two models, we strive to get beneficial properties from both. Our approach combines the safety of a reactive safe model and the performance from a predictive model.
@@
Figure 3 shows how our algorithmic security mechanism works. The whole process starts when a new CGM reading arrives. Once the BioKernel receives a new CGM reading, it collects a recent history of CGM readings and insulin dosing and sends these as inputs to both the reactive safe model and the predictive ML model. These models independently produce pump commands, called a temporary basal command. The temporary basal command sets the insulin delivery rate for the insulin pump for a 30-minute duration. The BioKernel will collect both temporary basal commands and look back at the differences between the reactive safe model and predictive ML model for the past three hours to see how much insulin delivery it should attribute to the predictive ML model. If the predictive ML model is still within its safe insulin delivery limits, it uses the predictive ML model’s temporary basal command to program the pump. If it has exhausted its limits, we clamp it to fit within the insulin bounds we set relative to the reactive safe model.
To set bounds, we base our calculations on human physiology. People produce a background level of glucose for energy. People living with T1D need to inject insulin to facilitate the absorption of this background glucose.
Our method for setting bounds allows the predictive model to use up to three hours worth of background insulin in advance. If the prediction is correct and glucose levels rise (e.g., from eating), the proactive insulin will manage it. If the prediction is wrong and no extra glucose appears, the body’s ongoing glucose production will help use the extra insulin over the next three hours.
In the worst case, an incorrect prediction might lead to low glucose. However, with our mechanism in place, we limit the impact by our three-hour bound and offset by the body’s continuous glucose production.

### v4

@@ Abstract.
GlucOS introduces a novel architecture that allows users to personalize diabetes management using any predictive model (including ML) for insulin dosing while simultaneously protecting them against malicious models. We also introduce a novel holistic security mechanism that adapts to unprecedented changes to human physiology. We use formal methods to prove correctness of critical components and incorporate humans as part of our defensive strategy. Our evaluation includes both a real-world deployment with seven individuals and results from simulation to show that our techniques generalize. We highlight that our results are not from a lab study, with people using GlucOS to manage Type 1 Diabetes in their daily lives. Our results show that GlucOS maintains safety and improves glucose control even under attack conditions. This work demonstrates the potential for secure, personalized, automated healthcare systems. Our entire source code is available at this link.
@@ 1. Introduction
Type 1 Diabetes (T1D) is a metabolic disorder where an individual’s pancreas stops producing insulin. To compensate, they inject synthetic insulin. Mobile applications, called automated insulin delivery systems, use subcutaneous sensors to continuously monitor glucose concentrations (glucose is the body’s primary energy source) and regulate glucose by automatically injecting insulin via an insulin pump (Figure 1). However, insulin is a dangerous hormone as too much insulin can kill people in a matter of hours (cryer2012severe, ) and too little insulin can kill people in a matter of days (benoit2018trends, ). For automated insulin delivery systems, the key challenge is to maintain a balance: provide enough insulin to prevent dangerously high glucose levels while avoiding excessive insulin that could lead to life-threateningly low glucose levels. In this context, security means ensuring integrity for insulin dosing against malicious attacks or inadvertent errors.
@@
Several commercial (beta_bionics, ; tandem_iq, ; omnipod, ) and open-source (loop, ; openaps, ) automated insulin delivery systems exist today, but none of them consider security as a primary design constraint. The complexity of these systems makes it difficult to reason about their correctness, which is crucial for security. For example, the OpenAPS (openaps, ) core function for calculating insulin doses consists of 1192 lines of Javascript code, 63 input and configuration parameters, and 90 branch statements. Furthermore, automated insulin delivery systems that use the OpenAPS algorithm embed it into native apps using a WebView for Javascript interpretation, increasing the overall complexity and attack surface. Vulnerabilities or attacks on such systems can be severe and fatal.
In this paper, we take on the challenge of building the first trustworthy automated insulin delivery system, called GlucOS. Our design includes: (1) an architecture where we apply separation principles for isolated and simple components, (2) a novel security mechanism and policy to enable secure insulin delivery, (3) a mechanism to account for unprecedented changes in human physiology, and (4) the application of formal methods to prove correct the implementation of our insulin delivery path, the most sensitive part of the system. Despite these efforts, we identify scenarios that require users to take action for security. Thus, we incorporate humans as part of our defensive strategy, providing them with agency and trust to take the required, corrective actions. In contrast to existing systems, GlucOS enables the use of any algorithm (including ML) to calculate insulin doses, while ensuring security by moderating insulin doses using 25 lines of formally verified Swift code.
Our approach is distinguished by its holistic consideration of the entire problem of automated insulin delivery. We provide clear security boundaries, mechanisms and policies to mitigate attacks on vulnerable components, and formal methods to provide assurances that our implementation is correct. Unlike prior research that exclusively focuses on simulation for evaluation (zhou2021data, ; zhou2023knowsafe, ), we also evaluate our approach with an end-to-end system used by individuals to manage diabetes with all the uncertainties of the real world.
Given their life-threatening stakes, automated insulin delivery systems require special consideration beyond the traditional purview of the systems and security communities. Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating their impact. Our position is that we need to make sure that automated insulin delivery systems are never assigned CVEs, because vulnerabilities can kill people. Thus, we focus on building a new system, using security-first principles, to avoid security vulnerabilities and mitigate their damage from the start.
@@
  • Security mechanisms and policies that protect individuals from malicious dosing algorithms, as well as drastic changes in human physiology.
@@ 2. Overview
Doctors and healthcare professional recognize Type-1 Diabetes as a predominantly self-managed condition and support the adoption of automated insulin delivery systems for management (pho, ; khor2023individualized, ; midyett2023one, ).
@@ 2.1. Locking down the insulin delivery path
Our system architecture focuses on providing security mechanisms and policies on the insulin delivery path. The insulin delivery path in automated insulin delivery systems includes models for calculating the amount of insulin still active in the individual (insulin can take up to six hours for complete absorption), closed-loop algorithms for calculating and delivering insulin doses that could use deep neural networks to predict future metabolic states. Together, these components are both the most complex, and the most sensitive from a security perspective because they are responsible for insulin delivery.
@@ 2.2. Architecture
The BioKernel handles the typical operations of automated insulin delivery systems. It runs the closed-loop algorithm every five minutes as the new CGM readings come in. Once the BioKernel receives new CGM readings, it collects the current metabolic state of the individual and sends it to a predictive model. The predictive model calculates the amount of insulin to inject based on its assessment of the individual’s current metabolic state and predictions of future metabolic states. The BioKernel then issues commands to the pump to deliver the calculated amount of insulin. A pump driver converts these commands into low-level I/O to program the pump and adjust insulin delivery.
To withstand attacks, the BioKernel serves as a reference monitor (anderson1972computer, ) to enforce security policies on the predictive model’s pump commands. This reference monitor architecture ensures that we can withstand attacks from malicious predictive models without having to know the internals of how they operate. Our reference monitor enforces algorithmic security, where we bound the amount of insulin injscted by the model to stay within dynamic safety bounds computed by a simple reactive safe model.
Our use of formal methods focuses on the software in the insulin delivery path. We use Hoare logic to prove key functions correct, define system states and transitions to handle runtime verification check failures safely, and we introduce the notion of a novel biological invariant. Our biological invariant is an end-to-end check of our theoretical calculations of how the individual’s metabolism should behave compared to what we observe in practice. Violations of this invariant suggest the need to temporarily disable automated insulin delivery in response to an unprecedented change to human physiology.
@@ 2.3. Threat model
Automated insulin delivery systems inject synthetic insulin to regulate glucose concentration in the body. In the long run, high glucose or hyperglycemia leads to dangerous health outcomes such as kidney failure, heart disease, amputation etc. The most serious and immediate health risk for people who use synthetic insulin is low glucose, or hypoglycemia. Hypoglycemia results from an overdose of insulin. Once an automated insulin delivery system has delivered insulin, there is no way to remove it from the body. If untreated, severe hypoglycemia can lead to loss of consciousness, seizure, or death in a matter of hours. Thus, our primary security objective is to reduce or eliminate hypoglycemia.
@@
Since personalization is important for T1D self-management (glucomaker, ), our design philosophy centers around humans. We provide flexibility for people to use any model for insulin dosing that suits their preferences. Our goal is to make sure that they stay safe even when they pick vulnerable or malicious models. Attackers may target automated insulin delivery systems to directly harm the human using the system.
For safety, we focus on protecting users from incorrect insulin doses that emerge from both, poorly chosen insulin dosing models, as well as inadvertent mistakes from benign models. Malicious or vulnerable models, if unchecked, can be devastating because they directly affect insulin delivery and can lead to life-threatening hypoglycemia.
We build GlucOS as a complete iOS app, and accordingly assume that the Swift type system, that forms the backbone of the iOS ecosystem, is secure. We only consider FDA-approved CGMs and pumps. With the FDA being stringent with security guidelines for medical devices (fda_medical, ), we assume that CGM and pump hardware are free of attacks. However, we assume that CGMs and pumps are prone to measurement errors within their standard error bounds (garg2022accuracy, ; dexcom_error, ; Zisser2011OmniPodAccuracy, ).
We discuss malicious pump drivers and ways to overcome them in Appendix D.
3. Design principles
Four key principles guide our design of GlucOS:
  • Principle 1: Focus on simplicity for security mechanisms by using domain knowledge. Rather than design a general security mechanism for a broad class of problems, we focus specifically on the highly-risky domain of automated insulin delivery. Using domain-specific insights, we provide simple security mechanisms that both humans and theorem proving software can easily reason through.
  • Principle 2: Design the system to empower humans to provide security. Automated insulin delivery represents an extreme example of high-stakes healthcare. For people to be able to use a secure system to manage their health, they need to have agency over the security of their system. Rather than simply “alerting the human” if something unexpected happens, we empower the humans using GlucOS to be active participants in the security of their system.
  • Principle 3: Consider asymmetries in risk. Insulin dosing possesses an inherent asymmetry where overdosing poses greater immediate danger than underdosing (cryer2012severe, ; benoit2018trends, ). With GlucOS, we optimize our mechanisms to be more aggressive when protecting individuals from insulin overdose.
  • Principle 4: Adaptability for security. Although we carefully design our software to be correct, humans and glucose metabolism can change quickly. Any change that affects the correctness of models of glucose metabolism would compromise the security of the entire system. Thus, we design GlucOS to adapt to constantly changing human physiology to provide security both in our formulation of the problem and our use of formal methods.
Although we do strive to apply these principles to the entire design, there are times when there is tension between them. For example, our biological invariant (Section 5) helps with adaptability (Principle 4) by detecting when our physiological models are wrong and considers asymmetries in risk (Principle 3). But to ensure that the human remains empowered (Principle 2), we use necessarily complex logic to deal with cases when the biological invariant does not hold (violation of Principle 1). These types of tensions are fundamental to a real-world system that manages human health and underscores the importance of building real systems to gain experience to navigate these tradeoffs.
4. Algorithmic security: Withstanding malicious dosing algorithms
We have dual goals for our algorithmic security mechanism of being both defensive and enabling. Our security mechanism is defensive by protecting people from inappropriate insulin doses emerging from bugs, corner cases, or even attacks merging from complex predictive algorithms – be it deep neural networks or massive sets of hand-crafted rules. Our algorithmic security mechanism is enabling in allowing people to employ any predictive algorithm to personalize the system to suit their individual preferences.
To enable predictive algorithms, GlucOS’s algorithmic security mechanism introduces a novel architecture that separates the use of predictive models for insulin dosing from accounting and safety logic to ensure that they are operating safely. This architecture provides flexibility in using any predictive model while enforcing bounds on insulin doses with a simple and formally verified model for safety.
4.1. Need for algorithmic security
Closed-loop automated insulin delivery systems have traditionally used reactive physiological models (such as PID controllers (weinzimer2008fully, ; kushner2018data, )) to calculate insulin doses. These models essentially use standard physiological models to react to offset the amount of excess glucose in the body. This makes it easy to reason about the correctness of their insulin doses. However, because synthetic insulin is slow-acting and takes up to six hours to fully absorb, the reactive nature of these models limits their effectiveness in regulating elevated glucose concentrations. Thus, having the ability to inject insulin proactively by predicting glucose concentrations leads to more effective glucose regulation. But, predictive dosing opens up the individual to the risk of potentially dosing too much insulin, and current insulin delivery systems have been reluctant to adopt the most advanced forms of predictive models.
@@ 4.1. Need for algorithmic security
Researchers have shown that modern advances in ML, such as deep neural networks, work well in making predictions for automated insulin delivery (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ). Despite their promise, none of the existing automated insulin delivery systems use deep neural networks. Instead, they use statistical methods, like linear regression or model predictive control (physio2011, ; physio2019, ; lunze2013control, ; control2019, ; openaps, ). The reason for using this more traditional form of ML is sound. It provides consistent, explainable results and has a physiological basis that people can use to vet its decisions. While advanced deep neural networks have more predictive power, current systems do not incorporate them due to their black-box nature and the potentially dire consequences of mispredictions. Even a well-tested ML model does not guarantee immunity to misprediction (nfl_ml, ; nfl_adversarial, ), especially when predicting human metabolism that is influenced by numerous ever-changing factors such as food, exercise, stress, environment, allergies, puberty, temperature and so on (fourtytwofactors, ). An automated insulin delivery system’s unchecked misprediction or “hallucination” (chatgpt_hallucination, ) could have lethal consequences if it delivers an inappropriate insulin dose (cryer2012severe, ; benoit2018trends, ).
Beyond the use of ML, the complexity of rule-based predictive models (such OpenAPS’ thousands of lines of code (openaps, )), make it difficult to reason about their correctness, even if they are generally effective at regulating glucose for diabetes management. Additionally, people requesting enhancements to such models (cob, ; hypo_protect, ) based on their personal struggles further exacerbates model complexity.
Thus, avoiding predictive models for insulin delivery comes at the cost of effective diabetes management. On the other hand, incorporating them introduces safety concerns stemming from challenges in the explainability of complex code and deep neural networks.
@@ 4.2. Mechanism fundamentals
The key novel insight underpinning our security mechanism design is that all correct insulin dosing algorithms, regardless of whether they are predictive or reactive, will deliver the same amount of insulin over a sufficiently long period (such as a full digestion-absorption cycle). People inject insulin to facilitate glucose absorption after eating. Digestion converts food to glucose, and insulin facilitates the absorption of glucose by tissues (e.g., the brain) where the body can use it as an energy source. Since insulin dosing is dominated by food consumption, all correct algorithms should dose the same amount of insulin to cover glucose from digestion. But the timing of these insulin doses has a large impact on the individual’s glucose regulation.
Our algorithmic security mechanism pairs two models. Our first model is a predictive model that anticipates future metabolic states and doses insulin proactively. Our second model is a reactive safe model that measures the current metabolic state and doses insulin based only on what it can measure currently. This pairing balances proactive insulin dosing for faster responses with a more conservative approach for safety. As per our insight, the reactive model would dose enough insulin eventually, so we use it as our baseline but provide predictive models with enough buffer to inject insulin earlier and safely.
Architecturally, in a novel approach for automated insulin delivery systems, the predictive model runs outside of our trusted computing base while the reactive safe model is part of the trusted computing base. This separation enables individuals to use whatever predictive models they want, while maintaining security. It also defends against malicious predictive models, as all insulin delivery decisions are ultimately vetted by the reactive safe model. This design also enables people to update their predictive models without violating core safety guarantees.
Conceptually, the predictive model is the primary model in our system and should control the pump most of the time. We use the reactive safe model for accounting so that we can track how far the predictive model deviates from a known-to-be-safe baseline. Then, we bound the size of this deviation. By combining these two models, we strive to get beneficial properties from both. Our approach combines the safety of a reactive model and the performance from a predictive model.
@@
Figure 3 shows how our algorithmic security mechanism works. The whole process starts when a new CGM reading arrives. Once the BioKernel receives a new CGM reading, it sends the recent history of CGM readings and insulin doses to both the reactive safe model and the predictive model. These models independently produce pump commands (called a temporary basal command) that set the insulin delivery rate for the pump over the next 30 minutes. The BioKernel then collects both commands and looks back at the differences in doses between both models over the past three hours to calculate how much insulin the predictive model has dosed. If the predictive model is still within the safe insulin delivery bounds relative to the reactive safe model’s doses, the BioKernel uses the predictive model’s command to program the pump. If the predictive model has exhausted its limits, we clamp its calculated insulin to fit within the insulin bounds relative to the reactive safe model.
To set bounds, we base our calculations on human physiology. People produce a background level of glucose for energy. Our method for setting bounds allows the predictive model to use up to three hours worth of background insulin in advance. If the prediction is correct and glucose levels rise (e.g., from eating), the proactive insulin will manage it. If the prediction is wrong and no extra glucose appears, the body’s ongoing glucose production will use the extra insulin over the next three hours.